the integral of f(x) = f(x)


then what is the integral of f(x)^2 ?

Grade:12

1 Answers

Prudhvi teja
13 years ago

Dear hsn

if integral of f(x) = f(x)

then f1(x) = f(x)

so integral of f(x)f1(x)dx = f(x)2/2

Please feel free to post as many doubts on our discussion forum as you can.we will get you the answer and detailed  solution very  quickly.

All the best.

Regards,
Askiitians Experts
Prudhvi Teja

 

 


Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian